# PertEval: Unveiling Real Knowledge Capacity of LLMs with Knowledge-Invariant Perturbations

### v1

@@ 3.3 Transition Analysis For Measuring Real Knowledge Capacity
Here $I(\cdot)$ denotes the indicator function.
Aspect 1: Overall Performance Stability. In terms of overall performance, we expect LLMs that have robustly acquired knowledge/skills required by benchmark datasets should have a stable performance when confronting knowledge-invariant perturbations. To this end, we propose Performance Drop Rate (PDR) to measure the over performance stability of LLMs given knowledge-invariant perturbations:
$\text{PDR}(M,D,\sigma)=\frac{1}{|D|}\sum_{x\in D}I[M(q_{\sigma(x)})=y_{\sigma(x)}]-I[M(q_{x})=y_{x}].$ (2)
Here $M$ denotes the LLM, $D$ and $D^{\prime}$ denotes the original and the perturbed dataset respectively. $PDR$ essentially equals to the discrepancy between the LLM’s accuracy on $D$ and that on $D^{\prime}$. If $PDR<0$, then the perturbation decreases the overall performance of LLMs. Then the LLM does not robustly acquire knowledge/skills. To further obtain reliable conclusions, we further make Wilcoxon signed-rank test for original and perturbed question sample pairs.
Table 2: Performance Transition Matrix
Number Correct after perturbation $\sigma$ Incorrect after perturbation $\sigma$
Correct before perturbation $\sigma$ $CC$ $IC$
Incorrect before perturbation $\sigma$ $IW$ $CW$
Aspect 2: Correct Response Consistency. To this end, we propose the Recall of Performance (ROP) metric. Its name originates from an analogy to the classical Recall score. To define ROP, we first propose the performance transition matrix, which is similar to the confusion matrix, as shown in Table 2. Here the first capital character of each variable (i.e., C/I) denotes “consistent” or “inconsistent”. The second capital character of each variable (i.e., C/N) denotes “correct” or “wrong”. For example, the variable $CC$ means the number of “consistent correct” before and after the perturbation. Next, similar to recall score, ROP is defined as the ratio of consistent correct in the number of all correct samples before perturbation, i.e.,
$\text{ROP}(M,D,\sigma)=\frac{CC}{CC+IC}.$ (3)
Tha range of ROP is in $[0,1]$. The higher the ROP, the better the correct response consistency given the perturbation. It should be noticed that, unlike PDR, there does not exist a “threshold” for ROP to measure the correct response consistency. Rather, the ROP serves as a score of correct response consistency, which can be utilized to compare the performance of different LLMs and strategies.
(a) Knowledge invariance score rated by GPT-4.
(b) Logarithm of Levenshtein distance.
Figure 1: LLM-based knowledge invariance scoring using systematic sampling (interval = 10).
@@ 4 Experiment
Datasets. In the experiment, to comprehensively cover evaluation domains and subjects while keeping the evaluation cost, we select test data in College Mathematics (C-Math), World History (W-History), Professional Psychology (P-Psychology) and Professional Medicine (P-Medicine) from the Massive Multitask Language Understanding (MMLU) benchmark, with each representing a supercategory in STEM, Humanities, Social Sciences and Other respectively. Statistics of selected datasets are presented in Table 9 in Appendix D.
@@
Large Language Models. Based on existing LLM leaderboards like OpenCampass [^1] https://rank.opencompass.org.cn/ and considering both the popularity and timeliness of LLMs, we select six representative LLMs for evaluation. These LLMs include close-sourced (gpt-4-turbo, gpt-3.5-turbo, gemini-1.0-pro, glm-3-turbo) and open-sourced (mistral-7b-instruct-v0.2, llama-3-8b-instruct).
4.1 Checking Knowledge Invariance for Perturbations
@@
Knowledge invariance checking results are presented in Figure 1 and Table 3. For LLM-based knowledge invariance scoring (Figure 1), we compare the score of proposed perturbations with the baseline PromptAttack[36] and find the excellent average score ($>4.5$) for each proposed perturbation. We also calculate the Levenshtein distance (edit distance) of perturbed and original questions for each perturbation, and we find that the traditional edit distance is not suitable for measuring knowledge invariance because of a lack of the correlation between edit distance and knowledge invariance score. For testing on mastered questions (Table 3), we find that most perturbations obtain an average PDR closed to zero, which indicates their consistent performance on the perturbed and the original data. The PDR of SwapPos on World History and Professional Psychology is less than -0.05 although it does not change any knowledge-relevant information based on our design. A possible explanation is that SwapPos changes the global ordering of question prompts, thus affects the text generation of these self-regression-based LLMs. More detailed result analyses are available at Appendix D.1
4.2 Real Knowledge Capacity Evaluation (Transition Analysis Part 1)
Figure 2: Real knowledge capacities measured by ACC@Consist with composite knowledge-invariant perturbation. Metrics are averaged on datasets. ACC@Original denotes accuracy from the original data. ACC@Perturb denotes accuracy from the perturbed data. ACC@Consist denotes ratio of consistent correct responses in both the original and the perturbed data.
@@ 4.2 Real Knowledge Capacity Evaluation (Transition Analysis Part 1)
In PertEval, each knowledge-invariant perturbation strategy focuses on a specific question transformation. To construct the most challenging test scenario for LLMs and obtain reliable and conprehensive evaluation results, we compose all knowledge-invariant perturbations together to obtain the composite perturbation to quantify LLMs’ real knowledge capacity. Then the real knowledge capacity of LLMs is measured by ACC@Consist. Evaluation results are shown in Figure 2. Based on these results, we obtain several findings in terms of LLMs’ knowledge capacity evaluation:
@@
1. The knowledge capacity of LLMs is significantly overvalued in static dataset-based benchmarks. Comparing the ACC@Original with ACC@Perturb and ACC@Consist for each LLM, we observe a significant drop of LLMs’ performance on perturbed dataset. Especially for gpt-4-turbo, gpt-3.5-turbo and gemini-1.0-pro, the change exceeds 20%. Even the most powerful selected LLM, gpt-4-turbo, gains only ACC@Consist as 0.5551. These results indicate a large gap between evaluated knowledge capacity of LLMs in static data-based benchmarks and real knowledge capacity of LLMs in complicated scenarios.
@@
2. Every selected LLM has its own consistently mastered knowledge. This result is deduced from a comparison between LLMs’ performance and pure guess. Given $k$ options and a single answer, the expected ACC@Consist is $1/k^{2}$. Therefore, the expected of ACC@Consist for pure guess is $0.06251/4^{2}=$ 0.0625 in the experiment ($k=4$)222see Appendix D.2 for detailed demonstration.. We observe from Figure 2 that ACC@Consist values of all selected LLMs are significantly larger than 0.0625, which means that each selected LLM indeed has its own consistently mastered knowledge, although most of them master less than a half (ACC@Consist < 0.5) of the total knowledge.
4.3 Response Pattern Analysis (Transiton Analysis Part 2)
(a) Original data
(b) Perturbed data with composite perturbation
@@ 4.3 Response Pattern Analysis (Transiton Analysis Part 2)
Figure 3: Response patterns of gpt-4-turbo in all datasets.
@@
We conduct response pattern analysis to explore how perturbations affect the performance of LLMs, and we discover that knowledge-invariant perturbations affect LLMs by increasing the ratio of extra multiple choices. As presented in Figure 3, the main reason why the performance of gpt-4-turbo drops dramatically is that they select extra incorrect options more frequently in the perturbed data (gpt-4-turbo: $12\%\to 21.7\%$). These LLMs tend to select extra incorrect options besides selecting correct options. A possible explanation of this phenomenon is that the LLMs memorize correct options, while lack the ability to recognize and filter out incorrect options, which means that these LLMs do not really master knowledge of the questions. Therefore, besides the ability to distinguish correct answers, we should also emphasize enhancing large language models’ ability to recognize and filter out incorrect answers when facing options in real-world problems. More detailed response pattern analysis results are available at Appendix D.3.
4.4 Overall Performance Stability (Transition Analysis Part 3)
@@
This analysis aim to unveil the vulnerability of LLMs to different perturbations in terms of overall performance. Evaluation results of overall performance stability are shown in Table 4. Here each value in the table’ cell is the macro PDR, i.e., the average value of PDRs separately calculated on each dataset. The Wicoxon signed-rank test is applied to the micro PDR, i.e., the overall PDR calculated on the all-in-one dataset. Specifically, let $s=I(M(q_{x})=y_{x})\in\{0,1\}$ and $s^{\prime}=I(M(q_{\sigma(x)})=y_{\sigma(x)})\in\{0,1\}$ respectively denotes model $M$’s score on the original and the perturbed questions, the alternative hypothesis of the Wilcoxon signed-rank test is $H_{a}:s>s^{\prime}$, or, $H_{a}:s^{\prime}-s<0$. From Table 4, we can both analyze for each column or for each row. For columns (strategy-aspect), we first observe that all selected LLMs shows negative macro PDRs given the content-level perturbation, the knowledge-invariant paraphrasing (KnInvPara). In addition, for close-sourced LLMs, the micro PDR is significantly negative, which means that the these LLMs significantly lack the content-level knowledge acquirements robustness in these datasets. Another observation is that all the LLMs are highly sensitive to the SwapPos perturbation. We suppose the reason is that, unlike other format-level perturbations that mostly focus on local format, SwapPos changes the global format of questions, which highly destroys self regression-based LLMs’ text generation process, even though this perturbation is entirely knowledge-invariant.
@@
glm-3-turbo -$0∗$0190 -0.1397∗∗ -0.0118 +0.0522 -0.2142∗∗ -0.0667
mistral-7b-instruct-v0.2 -0.0264 -$0$0.2789∗∗ +0.0793 -0.0844∗∗ -0.1275∗∗ -0.0763
llama-3-8b-instruct -0.0336∗∗ -0.0091 -0.0939∗∗ -0.0368∗∗ -0.2920∗∗ -0.1814∗∗ -0.1074
AVG -0.0456 -0.0087 -0.0994 0.0003 -0.0596 -0.1335
@@
glm-3-turbo 0.7164 0.6756 0.5943 0.7884 0.6816 0.4026 0.6288
mistral-7b-instruct-v0.2 0.6543 0.6202 0.1339 0.7679 0.5128 0.4060 0.5158
llama-3-8b-instruct 0.8055 0.7433 0.7541 0.8102 0.3971 0.5209 0.6676
AVG 0.7522 0.7624 0.6673 0.8323 0.6858 0.6009
4.5 Correct Response Consistency (Transition Analysis Part 4)
@@
This analysis aim to unveil the vulnerability of LLMs to different perturbations in terms of the consistency of correct responses. Experimental results of ROP are presented in Table 5. First, in terms of LLMs, the correct response consistency of different LLMs vary a lot. Among them, gpt-4-turbo performs best with the average macro ROP as 0.8949, which is a big gap compared to other LLMs. We also observe that gemini-1.0-pro performs well especially on the OptionForm and OptionCaesar strategies (Macro PDR > 0.9), which can detect the selection bias of LLMs. Without prerequisite information, we suppose that special effort have been made in the training of gemini-1.0-pro to overcome selection bias. For other LLMs, their correct response consistencies are vulnerable to almost all knowledge-invariant perturbations, which means that their correct answer in the original datasets are less reliable. On the other hand, the effect of different strategies on correct response consistency also vary a lot. Similar to results in overall performance stability, SwapPos is the most powerful strategy that influences LLMs’ correct response consistency, which exposes the inherent flaw of LLMs in defending global order-level perturbation.
5 Conclusion
In this paper, we proposed PertEval, an evaluation toolkit for measuring large language models’ real knowledge capacity using human-like knowledge-invariant perturbations and transition analysis methods. Based on the well-defined knowledge perturbation strategies and metrics in transition analysis, we revealed the overestimated knowledge capacity of six state-of-the-art LLMs on the MMLU benchmark. Response pattern analyses discovered that PertEval retains LLMs’ uncertainty to specious knowledge and revealed their rote memorization to correct options. Further transition analysis by PertEval illuminated the vulnerability of different LLMs to different perturbations and could guid the development of refinement. Finally, we discussed ways to enhance the knowledge capacity of LLMs using PertEval.

### v2

@@ 3.3 Response Consistency Analyses for Measuring Real Knowledge Capacity
Here $I(\cdot)$ is the binary indicator function. Taking a step further, we implement a response pattern analysis that uncovers the causes of the discrepancy between ACC@Consist and the original ACC.
Overall performance stability. In terms of performance stability, we expect that LLMs which have robustly acquired the knowledge and skills required by benchmarks should exhibit stable performance when faced with knowledge-invariant perturbations. To evaluate this, we propose Performance Drop Rate (PDR) as a metric to measure the overall performance stability of LLMs under knowledge-invariant perturbations.:
$\text{PDR}(M,D,\sigma)=\frac{1}{|D|}\sum_{x\in D}\Big{(}I[M(q_{\sigma(x)})=y_{\sigma(x)}]-I[M(q_{x})=y_{x}]\Big{)}.$ (2)
Here $D$ denotes the original test dataset and $\sigma$ is a perturbation strategy. Essentially, $PDR$ measures the discrepancy between the LLM’s accuracy on $D$ and the perturbed dataset. When $PDR<0$, the perturbation decreases the overall performance of an LLM, indicating that it does not robustly acquire knowledge and skills. To obtain more reliable conclusions, we further conduct Wilcoxon signed-rank test for original and perturbed question sample pairs.
Correct response consistency. We also adopt the Recall of Performance (ROP) metric to measure correct response consistency. The term “ROP” draws an analogy to the classical recall score. Let $CC$ denote the number of responses consistently correct before and after perturbation, and $IC$ the number of responses initially correct and later incorrect after perturbation (see Appendix B.3 for detailed demonstration). Similar to the recall score, ROP is defined as the ratio of consistent correct responses to the total number of correct responses before perturbation, i.e.,
$\text{ROP}(M,D,\sigma)=CC/(CC+IC)$ (3)
The value of ROP is within $[0,1]$, with higher ROP indicating better correct response consistency under perturbation $\sigma$. It should be noticed that, unlike PDR, there does not exist an absolute threshold for ROP to evaluate goodness. Rather, ROP serves as an index of correct response consistency, which can be utilized to compare the performance of different LLMs under perturbations.
Table 2: Knowledge invariance $\uparrow$ rated by human scorers. Four independent scores from different human scorer groups are presented in ascending order for each cell.
Method C-Math W-History P-Psychology P-Medicine
PromptAttack 2.3/2.4/3.0/3.9 2.2/2.2/2.3/2.8 1.3/2.4/2.8/2.8 1.6/2.5/3.5/3.6
PertEval (ours) 3.6/3.8/3.9/3.9 3.7/4.1/4.1/4.3 4.3/4.4/4.5/4.7 4.2/4.3/4.4/4.6
Table 3: Knowledge invariance $\uparrow$ rated by superior LLMs. Values (a/b/c) in each cell denotes the average knowledge invariance score rated by gpt-4-turbo, claude-3.5-sonnet, and llama-3.1-405b, respectively.
Method C-Math W-History P-Psychology P-Medicine
PromptAttack 3.2/3.6/3.6 3.2/3.3/3.7 3.9/3.9/3.7 4.1/4.3/4.2
PertEval (ours) 3.8/3.9/4.0 4.0/4.2/4.0 4.0/4.4/4.0 4.1/4.4/4.0
Figure 2: (left) Perturbation-wise knowledge invariance $\uparrow$ by gpt-4-turbo (systematic sampling, interval = 10); (right) edit distances between original question and perturbed questions.
@@ 4 Experiments
Datasets. We choose the test data in College Mathematics (C-Math), World History (W-History), Professional Psychology (P-Psychology) and Professional Medicine (P-Medicine) from the Massive Multitask Language Understanding (MMLU) [6] for evaluation. This is a trade-off between comprehensively covering evaluation domains/subjects and keeping the evaluation cost affordable. Note that each dataset represents a supercategory in STEM, Humanities, Social Sciences, and Other fields, respectively. The statistics of the selected datasets are detailed in Table 11 in Appendix C.1.
@@
Large Language Models. Based on leaderboards like OpenCompass [47] and considering both the popularity and timeliness of LLMs, we select six representative LLMs for evaluation. These LLMs include close-sourced (gpt-4-turbo [48], gpt-3.5-turbo [49], gemini-1.0-pro [50] and glm-3-turbo [51]) and open-sourced (mistral-7b-instruct-v0.2 [52] and llama-3-8b-instruct [53]) models.
4.1 Knowledge Invariance Verification for Perturbations
We first investigate the knowledge invariance property of our perturbations using the two methods developed in Section 3.2. Employing the LLM-based knowledge invariance scoring method, we compare the scores of the proposed perturbations with those by PromptAttack [41] as a baseline.
Experiment Setup. For knowledge invariance scoring, 10 samples at equal intervals for each subject are used for scoring. Since there are four subjects and two perturbation strategies (1 baseline + 1 PertEval) for each question, we have 10*4*2 = 80 question pairs in total. To ensure that each question pair have four independent scores, eight human scorers are engaged in human-based scoring. Human scorers are equally divided into four groups. Each group independently scores for all 80 question pairs. Therefore, each human scorer within a group scores for 40 shuffled samples. For LLM-based scoring, we choose gpt-4-turbo, claude-3.5-sonnet, and llama-3.1-405b as scorers. For fine-grained perturbation-wise scoring, we choose gpt-4-turbo as the scorer.
Result Analysis. Overall knowledge invariance scores are presented in Table 2 and 3. PertEval outperforms the baseline, and the score mostly exceeds 4.0, the borderline of knowledge-invariance. Scores of both PromptAttack and PertEval on C-Math do not exceed 4.0. This is because many samples in College Mathematics of MMLU are short mathematical reasoning questions that have many mathematical symbols and statements. Indeed, in many STEM subjects, requirements for knowledge capacity and reasoning ability often mix together in a test question. This points out considerable potential for future development in robust LLM evaluation in STEM subjects.
Next, the results presented in Figure 2 show that the average scores of our perturbations approach 5.0 (perfectly knowledge-invariant) in most cases, consistently outperforming those from PromptAttack. We also present the Levenshtein distance (i.e., edit distance) between the perturbed and original questions under each perturbation. It turns out that the traditional edit distance is not suitable for measuring knowledge invariance due to its lack of correlation with knowledge invariance.
@@ 4.1 Knowledge Invariance Verification for Perturbations
Within mastered question testing (results depicted in Table 4), we observe that most perturbations achieve an average PDR closed to zero, indicating consistent performance on the perturbed and the original data. The PDR of SwapPos on World History and Professional Psychology is less than -0.05 despite that it does not alter any knowledge-relevant information by design. A possible explanation is that SwapPos changes the global ordering of question prompts, thus affecting the token generation of these self-regressive LLMs. More detailed analyses of results could be found in Appendix D.
[…]
4.2 Real Knowledge Capacity Evaluation
@@
Each knowledge-invariant perturbation in PertEval targets a specific question restatement strategy. To construct the most challenging test scenario and obtain reliable and comprehensive evaluation results, we compose all perturbations to create a composite perturbation. The real knowledge capacity of LLMs is then quantified by ACC@Consist before and after the composite perturbation. The results, shown in Figure 3, yield several insights regarding the evaluation of LLMs’ knowledge capacity.
@@
1. Overvaluation on static benchmarks. The knowledge capacity of LLMs is significantly overvalued on static benchmarks. By comparing ACC@Original with ACC@Perturb and ACC@Consist for each LLM, we observe a notable performance drop when using the perturbed dataset. For instance, the performance of gpt-4-turbo, gpt-3.5-turbo, and gemini-1.0-pro decreases by more than 25%. Even the most capable LLM in our selection, gpt-4-turbo, achieves an ACC@Consist of only 0.555. These findings reveal a substantial gap between the evaluated knowledge capacity of LLMs on static datasets and their real knowledge capacity in complex scenarios.
@@
2. Robustly mastered knowledge. Each LLM does have robustly mastered certain knowledge. This conclusion is drawn by comparing the LLMs’ performance to pure guessing. With $k$ options and a single correst answer, the expected ACC@Consist for random guessing is $1/k^{2}$, which equals to 0.0625 for $k=4$ in our experiments (see Appendix D.1 for detailed demonstration). We observe from Figure 3 that the ACC@Consist values of all tested LLMs are significantly higher than 0.0625, indicating that each LLM has indeed mastered some knowledge consistently. However, most LLMs have mastered less than half of the total knowledge, as indicated by ACC@Consist values below 0.5.
@@
Figure 4: Response patterns of gpt-4-turbo. Left: original data; Right: perturbed data.
4.3 Response Pattern Analysis
@@
We further conduct a response pattern analysis to examine how perturbations affect the performance of LLMs. We find that knowledge-invariant perturbations affect LLMs by increasing the ratio of selecting extra incorrect options. As shown in Figure 4, the main reason for the significant performance drop in gpt-4-turbo is their increased frequency of selecting additional incorrect options on the perturbed data (i.e., $12\%\to 21.7\%$). These LLMs tend to select extra incorrect options alongside the correct ones. One possible explanation for this phenomenon is that the LLMs indeed lack the ability to distinguish uncertain knowledge and filter out distracting options, while they correctly answer these questions in the original dataset via rote memorization. In other words, these LLMs do not truly master the knowledge behind the questions. Therefore, it is crucial to not only enhance LLMs’ ability to identify correct answers but also bolster their capability to recognize and eliminate incorrect answers when confronted with distracting choices in real-world scenarios. More detailed response pattern analysis results can be found in Appendix D.2.
4.4 Overall Performance Stability
@@
Table 5 presents the overall performance stability results, in which each number is the macro-PDR, i.e., the average of PDRs separately calculated on each dataset. We also apply the Wicoxon signed-rank test to the micro-PDR, which is the PDR calculated across all datasets combined. Specifically, let $s=I(M(q_{x})=y_{x})\in\{0,1\}$ and $s^{\prime}=I(M(q_{\sigma(x)})=y_{\sigma(x)})\in\{0,1\}$ denote model $M$’s score on the original and perturbed questions, respectively. The alternative hypothesis of the Wilcoxon signed-rank test is $H_{a}:s>s^{\prime}$, or equivalently, $H_{a}:s^{\prime}-s<0$. We can analyze the results of Table 5 from both a column-wise and row-wise perspective. From the column-wise perspective, we first observe that all selected LLMs have negative macro-PDRs given the content-level perturbation, i.e., knowledge-invariant paraphrasing (KnInvPara). Additionally, the micro-PDR of close-sourced LLMs is significantly negative, meaning that these LLMs lack robustness in content-level knowledge acquisition for these datasets. Another prominent observation is that all LLMs are highly sensitive to the SwapPos perturbation. We hypothesize that this sensitivity is due to the global format change introduced by SwapPos, which disrupts the text generation process of self-regressive LLMs, even though the perturbation is entirely knowledge-invariant. From the row-wise perspective, we find that our perturbations cause performance drop in most cases, e.g., at least 0.04 and up to 0.1 on average, indicating the universal vulnerability of current generation LLMs.
@@
glm-3-turbo -$0∗$0190 -0.1397∗∗ -0.0118 +0.0522 -0.2142∗∗ -0.0616
mistral-7b-v0.2 -0.0264 -$0$0.2789∗∗ +0.0793 -0.0844∗∗ -0.1275∗∗ -0.0763
llama-3-8b -0.0336∗∗ -0.0091 -0.0939∗∗ -0.0368∗∗ -0.2920∗∗ -0.1814∗∗ -0.1078
AVG -0.0411 -0.0102 -0.1151 -0.0059 -0.1083 -0.1637
@@
glm-3-turbo 0.7164 0.6756 0.5943 0.7884 0.6816 0.4026 0.6432
mistral-7b-v0.2 0.6543 0.6202 0.1339 0.7679 0.5128 0.4060 0.5159
llama-3-8b 0.8055 0.7433 0.7541 0.8102 0.3971 0.5209 0.6719
AVG 0.7793 0.7530 0.6597 0.8340 0.6186 0.5544
4.5 Correct Response Consistency
@@
The results of correct response consistency using ROP are presented in Table 6. There is significant variation in ROP across different LLMs. Among them, gpt-4-turbo performs the best, with an average macro-ROP of 0.9, which is a substantial advantage over other LLMs. Gemini-1.0-pro also performs well, particularly with the OptionForm and OptionCaesar strategies (both having a Macro-ROP > 0.9). Other LLMs, however, are vulnerable to almost all knowledge-invariant perturbations, which means that their performance on the original datasets are less reliable. On the other hand, the effect of different strategies on correct response consistency varies considerably. Similar to the results for performance stability, SwapPos is the most influential strategy affecting LLMs’ correct response consistency, which exposes the inherent flaw of LLMs in defending global order perturbation.
5 Discussion
Conclusion. The trustworthiness of capacity evaluation is, and will always be one of the most significant issues in the development of LLMs. In pursuit of this goal, we proposed an evaluation toolkit, PertEval, for unveiling LLMs’ real knowledge capacity on close-ended benchmarks. We discovered not only a substantial overestimation of the knowledge capacity of LLMs by static benchmarks, but also LLMs’ uncertainty to distinguish specious knowledge without rote memorization. Detailed response consistency analyses revealed the vulnerability of LLMs to various knowledge-invariant perturbations, especially the content-level KnInvPara and the format-level SwapPos, which could inspire further refinement for LLMs’ knowledge mastery. Indeed, the ultimate goal of this study is not only trustworthy evaluation, but to propel the development of LLMs’ knowledge capacity. Our primary experiments have verified the feasibility of using PertEval-generated data to improve LLM robustness to various test conditions via supervised fine-tuning (see Appendix D.3). We believe that PertEval marks a significant step toward more trustworthy LLM evaluation.
Limitations & Future Work. The first limitation of this work is its scope regarding suitable benchmarks. PertEval is suitable for close-ended benchmarks and cannot be directly applied to open-ended ones. The challenges stem from the variety of open-ended questions and the difficulty of their objective evaluation. Future research could explore how the concept of knowledge-invariant perturbation can be adapted to improve the trustworthiness of open-ended evaluations. The second limitation of PertEval is the lack of controllability, a common issue in LLM-based text generation. Due to the complexity of LLMs, it is challenging to precisely control specific characteristics in paraphrased texts while keeping knowledge-invariance. In the future, we aim to achieve more fine-grained controllability in PertEval, such as quantitatively adjusting paraphrasing tone and length, to obtain more continuous and nuanced evaluation results of LLMs. Besides, all conclusions of this study are based on experiments conducted on a subset of MMLU due to budget constraints. We are seeking economic methods to expand our experiments and further substantiate our findings. Finally, the risk of adaptive optimization against PertEval exists as it is an open-sourced toolkit. However, the diversity of perturbations and the cost of adaptive optimization could effectively reduce the influence of adaptive optimization. By constructing a PertEval open evaluation platform, we can regularly develop and change perturbation strategies to prevent this procedure.
Acknowledgements
This research was supported by grants from the Joint Research Project of the Science and Technology Innovation Community in Yangtze River Delta (No. 2023CSJZN0200), the National Natural Science Foundation of China (No. 62337001), the Fundamental Research Funds for the Central Universities, and the Alibaba Research Intern Program. We thank all the volunteers for their massive efforts in supporting our experiments, including Yi Cheng, Xiaowen Zhang, and Anyu Chen at Alibaba Cloud Computing, Dingchu Zhang at Nanjing University, Yuyang Xu at Zhejiang University, and Yangyang Wang at Shandong University.